What is the measure of an interior angle of a regular polygon with 16 sides?
Answer by palanisamy(496)

In any polygon the sum of the exterior angles is 360.
So an exterior angle of a polygon with 16 sides = 360/16
= 22.5
Therefore the measure of an interior angle of a regular polygon
with 16 sides = 180-22.5 = 157.5 degrees